### v1

Abstract
Federated Learning (FL) enables collaborative model training across multiple devices while preserving data privacy. However, it remains susceptible to backdoor attacks, where malicious participants can compromise the global model. Existing defence methods are limited by strict assumptions on data heterogeneity (Non-Independent and Identically Distributed data) and the proportion of malicious clients, reducing their practicality and effectiveness. To overcome these limitations, we propose Robust Knowledge Distillation (RKD), a novel defence mechanism that enhances model integrity without relying on restrictive assumptions. RKD integrates clustering and model selection techniques to identify and filter out malicious updates, forming a reliable ensemble of models. It then employs knowledge distillation to transfer the collective insights from this ensemble to a global model. Extensive evaluations demonstrate that RKD effectively mitigates backdoor threats while maintaining high model performance, outperforming current state-of-the-art defence methods across various scenarios.
Index Terms:
Federated Learning, Defence Method, Backdoor Attacks
I Introduction
Federated Learning (FL) has emerged as a powerful paradigm that enables multiple clients to collaboratively train a shared global model without exchanging their private local data, thereby enhancing data privacy and security [14]. In FL, each client trains a local model on its own dataset and shares only the model updates with a central server, which aggregates these updates to form the global model. Despite its advantages, FL faces significant challenges that threaten the effectiveness and integrity of the resulting global model.
One critical challenge is the vulnerability of FL to security threats, particularly backdoor attacks. In these attacks, malicious clients introduce harmful updates into the training process by embedding hidden triggers within their model updates [15]. These hidden triggers are specific patterns or inputs that, when encountered by the global model, cause it to produce incorrect or malicious outputs. Malicious clients achieve this by manipulating their local training data and altering their model parameters, ensuring that the backdoor behaviour is incorporated into the global model during aggregation without being detected. Since these triggers are often designed to be stealthy inputs, and malicious clients train on both clean input data and inputs containing the trigger, the attacks can go unnoticed during normal validation. This stealthiness makes backdoor attacks particularly effective and dangerous threats to FL systems [25]. As a result, the compromised global model behaves normally on standard inputs but can cause significant harm when the triggers are activated, posing serious risks to the integrity and reliability of FL deployments.
Another major challenge is handling Non-Independent and Identically Distributed (Non-IID) data across clients. In real-world scenarios, clients often possess data that are not representative of the overall data distribution, leading to inconsistent global model performance and reduced convergence rates. This heterogeneity can cause the global model to perform poorly on some clients’ local data, undermining the benefits of collaborative learning [24]. The variations in local data distributions can result in conflicting model updates from different clients, making it difficult for the global model to generalize well across all participants.
Existing defence methods against backdoor attacks employ various strategies but are often limited by strict assumptions regarding data heterogeneity. Specifically, they assume data are Independent and Identically Distributed (IID) across clients and the proportion of malicious clients is relatively low. Robust aggregation methods, such as RLR [17] and FoolsGold [8], adjust the aggregation process to mitigate the influence of malicious updates but may struggle against sophisticated attacks and are less effective under Non-IID data distributions. Root data validation approaches, like FLTrust [4] and Baffle [2], attempt to validate updates using trusted data but can be circumvented by stealthy attacks that mimic benign behaviour. Differential Privacy methods, including FLAME [16], introduce noise and apply weight clipping to protect privacy and integrity but require precise tuning and often involve trade-offs that degrade model performance. Knowledge distillation-based methods, such as FedDF [13] and FedBE [5], address issues of Non-IID data and noisy gradients but may lack mechanisms for outlier detection, potentially allowing backdoor triggers to be transferred to the global model.
To overcome these limitations, we propose Robust Knowledge Distillation (RKD), a novel defense mechanism that enhances model integrity without relying on restrictive assumptions. RKD integrates clustering and model selection techniques to identify and filter out malicious updates, forming a reliable ensemble of models for knowledge distillation. Specifically, RKD employs cosine similarity and Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) to cluster client models and isolate potential outliers. Then, it selects models near the median of the benign cluster further to mitigate the influence of any remaining subtle malicious updates, thereby forming a trustworthy ensemble. Finally, it employs knowledge distillation to integrate the information from the ensemble into the global model, primarily reflecting benign update behaviour and mitigating subtle backdoor attacks, while also addressing Non-IID settings.
Our contributions are summarized as follows:
  • We introduce the RKD framework, a robust aggregation method for FL that counters sophisticated backdoor attacks without relying on restrictive assumptions about data distribution or the number of malicious clients.
  • We conduct extensive evaluations of RKD on diverse datasets, including CIFAR-10, EMNIST, and Fashion-MNIST, demonstrating that RKD effectively mitigates various backdoor attacks such as Adversarially Adaptive Backdoor Attacks to Federated Learning (A3FL) [22], Focused-Flip Federated Backdoor Attack (F3BA) [7], Distributed Backdoor Attack (DBA) [20], and Train and Scale Backdoor Attack (TSBA) [3].
  • Our empirical analysis shows that RKD outperforms baseline defense methods, particularly in scenarios with high poisoning rates and varying levels of Non-IID data. RKD maintains high model accuracy above $80\%$ while reducing attack success rates to below $17\%$, underscoring its superior capability in maintaining FL resilience.
Extensive experiments underscore the effectiveness of RKD in enhancing FL robustness, and maintaining high model performance while effectively mitigating backdoor threats. The implementation of the RKD framework is available at https://github.com/AnonymeSc/RKD.
II Background and Problem Setting
FL is a distributed machine learning paradigm where multiple clients collaboratively train a shared global model $M_{\text{global}}$ while keeping their local data $D_{i}$ decentralized and private. Each client $i$ trains a local model $M_{i}$ on its own dataset $D_{i}$ and then shares only the model parameters $\theta_{i}$ with a central server. The central server aggregates these local parameters to update the global model. A common aggregation algorithm is Federated Averaging (FedAvg), which computes the global model parameters as: $\theta_{\text{global}}=\frac{1}{N}\sum_{i=1}^{N}\theta_{i},$ where $N$ is the number of participating clients [14]. After aggregation, the server sends the updated global model $M_{\text{global}}$ back to all participating clients. Each client then updates its local model using the received global parameters, ensuring synchronization across the network. This collaborative process allows the construction of a global model that benefits all clients without exposing datasets, thus preserving data privacy [25].
However, despite its advantages, FL is vulnerable to security threats, particularly backdoor attacks. In backdoor attacks, malicious clients aim to compromise the integrity of the global model $M_{\text{global}}$ by manipulating their local training data $D_{i}$ and/or altering their model updates $\theta_{i}$ before sending them to the central server. These manipulations are designed to embed hidden malicious behaviours into the global model without being detected. Furthermore, the presence of Non-IID data distributions among clients exacerbates the difficulty in detecting and mitigating such attacks, as data heterogeneity can mask the malicious updates, making it challenging to distinguish between benign model variations and manipulations [7].
II-A Backdoor Attacks
II-A1 Data Poisoning Attacks
In backdoor data poisoning attacks, malicious clients intentionally modify their local training data to embed hidden triggers that induce the global model to exhibit malicious behaviour when these triggers are encountered [15]. Specifically, a malicious client takes its original dataset $D_{i}=\{(x_{j},y_{j})\}$ and creates a poisoned dataset $D_{i}^{\text{poison}}$ by injecting a small number of modified data instances $\{(x^{\prime}_{k},y^{\prime}_{k})\}$. In these instances, $x^{\prime}_{k}$ contains a trigger pattern—such as a specific pixel arrangement in an image—and $y^{\prime}_{k}$ is an incorrect target label chosen by the attacker. The modified dataset $\tilde{D}_{i}=D_{i}\cup D_{i}^{\text{poison}}$ is then used to train the local model $M_{i}$ [20].
The attacker’s objective is to train $M_{i}$ such that it performs well on normal data while misclassifying inputs containing the trigger pattern. This can be formalized by minimizing the loss function: $\mathcal{L}(M_{i}(\tilde{D}_{i}),y),$ where $y$ represents the labels. By successfully training on $\tilde{D}_{i}$, the model learns the backdoor association between the trigger and the target label $y^{\prime}_{k}$. When these malicious model updates $\theta_{i}$ are aggregated by the server, the global model $M_{\text{global}}$ inherits the backdoor behaviour, causing it to misclassify inputs containing the trigger pattern while maintaining normal performance on clean data.
II-A2 Model Poisoning Attacks
In backdoor model poisoning attacks, malicious clients manipulate their local model parameters $\theta_{i}$ to embed backdoors into the global model, often in combination with injecting triggers into their training data. Unlike data poisoning attacks that focus primarily on the training data, model poisoning attacks involve directly altering the model parameters to maximize the impact of the backdoor while minimizing the risk of detection [24].
In general, malicious clients may:
  • Manipulate Model Parameters: After training on the poisoned data, attackers further manipulate their model parameters $\theta_{i}$ to enhance the backdoor effect or to avoid detection. This can involve:
    • Adversarial Adaptation: Adapting the trigger and model parameters to remain effective against changes in the global model during training [22].
    • Scaling Weights: Amplifying the impact of the malicious update by scaling the model parameters [3]: $\theta_{i}^{\text{poison}}=\alpha\cdot\theta_{i},$ where $\alpha>1$ increases the update’s influence on the global model during aggregation.
    • Adding Perturbations: Introducing subtle changes to the model parameters to maintain stealth [7]: $\theta_{i}^{\text{poison}}=\theta_{i}+\delta,$ where $\delta$ is a small perturbation designed to embed the backdoor while avoiding detection by defence mechanisms.
Advanced optimization techniques, such as Projected Gradient Descent (PGD) [22], can be employed to iteratively adjust $\delta$ to find the optimal perturbation that balances attack success with stealthiness [19]. In some cases, attackers perform model replacement attacks, where they entirely replace their local model with a maliciously crafted model $\theta_{i}^{\text{poison}}$ and scale it to influence the global aggregation [3] disproportionately. By carefully manipulating both the training data and the model parameters, attackers ensure that their malicious updates blend in with those from benign clients, making the backdoor attack stealthy and difficult to detect.
II-B Impact on the Global Model
When the central server aggregates the model updates, the presence of malicious updates from attackers influences the global model parameters. Specifically, the aggregated global model parameters become: $\theta_{\text{global}}^{\text{poison}}=\frac{1}{N}\left(\sum_{i\in\mathcal{H}}\theta_{i}+\sum_{j\in\mathcal{M}}\theta_{j}^{\text{poison}}\right),$ where $\mathcal{H}$ is the set of honest clients and $\mathcal{M}$ is the set of malicious clients. The malicious updates $\theta_{j}^{\text{poison}}$ are designed to inject the backdoor into the global model $M_{\text{global}}^{\text{poison}}$. As a result, the compromised global model may behaves normally on standard (clean) inputs, maintaining high performance and thus not raising suspicion. However, when presented with inputs containing the backdoor trigger, the model exhibits incorrect or malicious behaviour, such as misclassifying the input or producing outputs desired by the attacker. This stealthy alteration poses significant risks to the integrity and reliability of the FL system, as the backdoor can remain undetected until the trigger is activated.
II-C Characteristics of Backdoor Attacks
Backdoor attacks often exhibit distinctive characteristics in the model updates sent by malicious clients, which can be exploited for detection. Key characteristics include:
  • Angular Deviation: Malicious model updates may have a different direction in the parameter space compared to updates from benign clients [3]. This directional difference can be quantified using the cosine similarity (or angular deviation) between the parameter vectors of malicious updates $\theta_{\text{attack}}$ and benign updates $\theta_{\text{benign}}$: $\Delta\theta_{\text{angular}}=\cos^{-1}\left(\frac{\theta_{\text{attack}}\cdot\theta_{\text{benign}}}{\|\theta_{\text{attack}}\|\|\theta_{\text{benign}}\|}\right)$. Angular deviations indicate that the malicious update is contributing in a different direction, which may be due to the attacker’s attempt to embed the backdoor like F3BA attack [7].
  • Magnitude Deviation: Malicious updates may have a significantly different norm magnitude compared to benign updates. This occurs when attackers scale their model updates to increase their influence on the aggregated global model [20]. The magnitude deviation can be observed when: $\|\theta_{\text{attack}}\|\gg\|\theta_{\text{benign}}\|.$
  • Subtle Deviations: Some attackers design their updates to closely resemble those of benign clients, keeping the deviation between the malicious and benign updates within a small threshold $\epsilon$ to avoid detection: $\|\theta_{\text{attack}}-\theta_{\text{benign}}\|<\epsilon.$ Techniques like A3FL attack [22] adjust both the magnitude and direction of the malicious update to embed the backdoor while maintaining stealthiness.
Understanding these characteristics is crucial for designing defence mechanisms that can detect and mitigate backdoor attacks without relying on overly restrictive assumptions.
II-D Threat Model
Our work assumes the FL system in which a subset of clients, referred to as malicious clients, aim to compromise the global model through backdoor attacks. The goal of these attacks is to embed hidden triggers into the global model, causing targeted misclassifications on specific inputs while maintaining high accuracy on clean data to avoid detection. Below, we outline the adversary’s knowledge, capabilities, and the defender’s knowledge:
  • Adversary’s Knowledge: The malicious clients have no access to other clients’ model updates or data. They operate under the assumption that they can only manipulate their local data and model updates.
  • Adversary’s Capabilities: The adversaries can collude and coordinate their attacks but cannot intercept or alter communications between other clients and the server. They can manipulate their local training data (data poisoning) and modify their model updates before sending them to the server (model poisoning).
  • Defender’s Knowledge: The server is aware that some clients may be malicious but does not know their identities. The server has access only to the submitted model updates and any auxiliary data (e.g., public unlabeled datasets) used for knowledge distillation.
III Related Work
Defending against backdoor attacks in FL is a critical area of research, with various strategies proposed to enhance the robustness of FL. These defence mechanisms can be broadly categorized into robust aggregation methods, clustering-based defences, and knowledge distillation approaches.
Robust Aggregation Methods aim to mitigate the influence of malicious updates during the model aggregation process. The Robust Learning Rate (RLR) method [17] adjusts the learning rates of clients based on the alignment of their updates with the global model’s direction. By assigning smaller learning rates to updates that deviate significantly, RLR reduces the impact of potentially malicious updates. However, RLR may not fully address attacks that manipulate the magnitude of updates, such as scaling attacks, and its effectiveness diminishes in heterogeneous (Non-IID) data environments where benign updates naturally vary in direction and magnitude.
FoolsGold [8] is designed to counter backdoor attacks by analyzing the similarity of gradient updates among clients. It assigns lower aggregation weights to clients whose updates are overly similar, under the assumption that malicious clients will produce highly similar gradients due to coordinated attacks. While effective against certain types of attacks, FoolsGold may inadvertently penalize benign clients with similar data distributions, leading to unfairness and potential degradation of overall model performance.
Clustering-Based Defences have been employed to distinguish between benign and malicious updates by grouping similar updates and identifying outlier models. Methods like FLAME [16] leverage HDBSCAN for clustering and incorporate noise addition to bolster security, yet they require improved adaptability to evolving threats. RFCL [1] also uses HDBSCAN to cluster clients and defend against gradient poisoning attacks but is less effective in defending against adaptive backdoor attacks [22]. A significant challenge with clustering-based defences is their reliance on the precision of the clustering process, which can be compromised in high-dimensional parameter spaces typical of deep learning models. Traditional clustering algorithms often struggle with high-dimensional data due to the curse of dimensionality, leading to less efficient clustering results and potential misclassification of benign updates as malicious or vice versa.
Knowledge Distillation Approaches have been recognized for their ability to improve learning efficiency and address Non-IID data distributions in a federated framework. Methods such as FedDF [13] and FedBE [5] leverage knowledge distillation to aggregate models. FedDF performs model fusion by distilling knowledge from client models into a global model using unlabeled public data, effectively aligning the models’ outputs. FedBE builds upon this by employing Bayesian ensemble techniques to instruct a student model based on the ensemble’s predictive distribution.
FedRAD [18] enhances FedDF by assigning weights to client models according to their median scores, which measure how often a model’s prediction corresponds to the median prediction among all clients. While these methods improve performance in Non-IID settings, they are vulnerable to backdoor attacks when transferring knowledge from ensemble models without analyzing outliers. Malicious clients can introduce backdoors into their models, and without mechanisms to detect and exclude these compromised models, the backdoor triggers can be propagated to the global model during distillation.
RKD builds upon existing methods by innovatively integrating and enhancing foundational techniques such as HDBSCAN and knowledge distillation. This approach addresses the limitations of prior works, particularly in defending against recent backdoor attacks in Non-IID data distributions.
First, by integrating cosine similarity with HDBSCAN, RKD improves clustering efficiency in high-dimensional parameter spaces, effectively identifying malicious updates even when they are subtle or adaptive. Focusing on the angular relationships between model updates captures essential differences without being overwhelmed by the volume of parameters, thereby ensuring accurate clustering despite Non-IID data.
Second, the RKD framework incorporates a robust model selection process that selects models near the median of the cluster. By identifying and using the most representative models, RKD further mitigates the influence of any remaining malicious updates. This selection forms a reliable and trustworthy ensemble for knowledge distillation.
Third, RKD ensures secure knowledge transfer by distilling from the carefully selected ensemble models, preventing the propagation of backdoor triggers to the global model. By excluding outlier models identified during clustering, RKD reduces the risk of incorporating malicious behaviours into the model. Additionally, knowledge distillation aids in smoothing out variations caused by Non-IID data, leading to a more generalized and robust global model.
By addressing the challenges of high-dimensional data, improving clustering efficiency, and ensuring secure knowledge transfer, RKD provides a robust defence against backdoor attacks in federated learning while effectively handling Non-IID data distributions.
IV Methodology
The RKD framework is designed to secure FL against backdoor attacks by identifying and mitigating malicious model updates. It consists of three core components: Automated Clustering, Model Selection, and a Knowledge Distillation Module. These components work together to detect and eliminate malicious influences while preserving the performance and integrity of the global model.
IV-A Framework Overview
The RKD framework employs a multi-tiered strategy to enhance model robustness within FL. This subsequent provides an overview of the RKD framework, outlining its key components and processes. Detailed explanations of each strategy method are presented in the subsequent subsections.
Initially, the central server initializes the global model $M_{\text{global}}^{(0)}$ and broadcasts it to all participating clients. Each client $i$ trains its local model $M_{i}^{(r)}$ on its private dataset $D_{i}$, starting from the current global model $M_{\text{global}}^{(r)}$. The result is updated model parameters $\theta_{i}^{(r)}$, which are then returned to the server.
At the server, the focus is on identifying potential malicious updates. The server computes the cosine similarity between each client model update $\theta_{i}^{(r)}$ and the current global model $M_{\text{global}}^{(r)}$. This metric captures the angular similarity between local updates and the global model, enabling the detection of updates that significantly deviate in direction—a characteristic of malicious behaviour known as Angular Deviation.
Using these similarity scores, the server employs the HDBSCAN algorithm to cluster the models based on their similarity to the global model. This clustering process distinguishes benign updates, which form dense clusters due to their similarity, from potentially malicious updates, which appear as outliers because of their dissimilarity. By clustering based on similarity scores rather than directly using high-dimensional parameter vectors, the computational complexity is significantly reduced, and the clustering becomes more scalable and efficient.
Within the benign cluster, the server computes the median of the model parameters by taking the median of each parameter across the models. This helps mitigate the impact of extreme values in the updates, addressing the Magnitude Deviation characteristic of malicious behaviour.
To further enhance robustness, the server selects models whose parameters are closest to the median to form an ensemble. This selection filters out any updates that may have evaded detection during clustering but still deviate from the central tendency of the benign models.
The selected models are then aggregated to form an initial distilled model. To further refine this model and improve its resistance to subtle backdoor triggers, while enhancing generalization across Non-IID data distributions, Knowledge Distillation (KD) is applied. The ensemble of selected models guides the refinement of the distilled model $M_{\text{global}}^{(r+1)}$, ensuring it replicates the collective behaviour of the ensemble models. This process helps protect against Subtle Deviations, where attackers craft updates that closely mimic benign updates in both magnitude and direction, making them harder to detect. Additionally, it improves performance across Non-IID.
Finally, the server broadcasts the refined global model $M_{\text{global}}^{(r+1)}$ to the identified benign clients for further training. By withholding the updated model from adversarial clients, RKD prevents attackers from adapting their strategies based on the latest changes, thereby reducing their ability to exploit subtle deviations. Clients classified as malicious continue to train using the previous global model $M_{\text{global}}^{(r)}$.
To further enhance security and mitigate the risk of adaptive attacks, RKD proposes an alternative solution. Instead of withholding the model entirely, the server sends a Perturbed Global Model (PGM) to clients suspected of being malicious This approach reduces the risk of adaptive attacks by obscuring the exact state of the global model while maintaining inclusivity, preventing attackers from deducing their classification. Malicious clients thus continue to train using this perturbed model, limiting their ability to exploit subtle deviations without providing them with precise information about the global model’s updates.
The perturbation added to the global model is intentionally kept minimal (e.g., noise scale of $1\times 10^{-4}$). This ensures that even if a benign client is mistakenly classified as malicious due to the non-IID nature of its data, the divergence from the true global model remains limited. The methodology allows for clients to be reclassified as benign in subsequent rounds. Once reclassified, they receive the unperturbed, updated global model, helping them realign with the overall training process.
Re-including previously misclassified benign clients may introduce slight variations in the global model’s performance. However, by keeping the perturbation scale minimal and allowing for dynamic reclassification, RKD mitigates this issue. The potential minor noise introduced is a reasonable trade-off for enhanced security against adaptive attacks.
Algorithm 1 summarizes the RKD framework.
Algorithm 1 RKD Framework Methodology
0:  Clients $\mathcal{A}$, number of iterations $R$
0:  Final global model $M_{\text{global}}^{(R)}$
1:  Initialize global model $M_{\text{global}}^{(0)}$
2:  for  $r=0$ to $R-1$  do
3:     if  $r==0$  then
4:        Send $M_{\text{global}}^{(0)}$ to all clients in $\mathcal{A}$
5:     else
6:        Send $M_{\text{global}}^{(r)}$ to benign clients $\mathcal{A}_{\text{benign}}^{(r-1)}$
7:     end if
8:     Collect models ${M}^{(r)}=\{M_{i}^{(r)}\mid i\in\mathcal{A}\}$
9:     Identify ${M}_{\text{benign}}^{(r)}$ and $\mathcal{A}_{\text{benign}}^{(r)}$ via clustering $\triangleright$ See Algorithm 2
10:     Select ensemble models $\mathcal{E}^{(r)}$ from ${M}_{\text{benign}}^{(r)}$
11:     Compute aggregated model $M_{\text{distill}}^{(r)}$ from $\mathcal{E}^{(r)}$
12:     Update $M_{\text{global}}^{(r+1)}=\text{KD}(M_{\text{distill}}^{(r)},\mathcal{E}^{(r)})$triangleright See Algorithm 3
13:  end for
14:  return  Final global model $M_{\text{global}}^{(R)}$
Automated Clustering. This component identifies and excludes potentially malicious models to maintain the integrity of the federated learning process. The method leverages cosine similarity and the HDBSCAN clustering algorithm to distinguish between benign and malicious model updates.
At iteration $r$, each client $i$ has local model parameters $\theta_{i}^{(r)}$. To identify alignment with the current global model, the server computes the cosine similarity $s_{i}$ between each client’s model and the global model $M_{\text{global}}^{(r)}$, which serves as the reference point: $s_{i}=\frac{\left(\theta_{i}^{(r)}\right)^{\top}M_{\text{global}}^{(r)}}{\left\|\theta_{i}^{(r)}\right\|\left\|M_{\text{global}}^{(r)}\right\|},\quad\text{for }i=1,\ldots,N.$
Here, $M_{\text{global}}^{(r)}$ represents the current aggregated global model. Higher cosine similarity values indicate greater alignment with the global model, which is expected for benign clients, while malicious clients are more likely to deviate.
These similarity scores $\{s_{i}\}$ are then clustered using HDBSCAN. Clustering based on scalar similarity scores significantly reduces computational complexity compared to clustering in the high-dimensional parameter space.
[…]
This dynamic adjustment ensures that the clustering remains adaptable to changes in client participation and training dynamics. For example, as the training progresses and fewer clients actively participate, $Q$ decreases, maintaining meaningful cluster sizes and computational efficiency.
Applying HDBSCAN to the set of similarity scores yields cluster labels $\{L_{i}\}$: $\{L_{i}\}=\text{HDBSCAN}\left(\{s_{i}\},Q\right).$
For each identified cluster $C_{k}$, the mean cosine similarity $\mu_{k}$ is computed: $\mu_{k}=\frac{1}{\left\|C_{k}\right\|}\sum_{i\in C_{k}}s_{i},$ where $\left\|C_{k}\right\|$ is the number of clients in cluster $C_{k}$.
The cluster with the highest mean cosine similarity $\mu_{\text{max}}$ is considered the benign cluster: $\mu_{\text{max}}=\max\left(\{\mu_{k}\}\right).$ Clusters are classified as follows:
Cluster $\text{Cluster }C_{k}\text{ is classified as }\begin{cases}{M}_{\text{benign}},&\text{if }\mu_{k}=\mu_{\text{max}},\\ {M}_{\text{malicious}},&\text{otherwise}.\end{cases}$
Models in the benign cluster are used for further training and receive the updated global model in the next iteration. Models in malicious clusters are excluded from receiving the updated global model, limiting their influence on the federated learning process. The detailed steps of the automated clustering procedure are in Algorithm 2.
Algorithm 2 Automated Clustering Algorithm
0:  ${M}^{(r)}=\{\theta_{i}^{(r)}\}_{i=1}^{N}$: Client models at iteration $r$
0:  ${M}_{\text{benign}}^{(r)}$, $\mathcal{A}_{\text{benign}}^{(r)}$: Benign models and client indices
1:  for each client $i=1$ to $N$ do
2:     Compute cosine similarity score $s_{i}=\frac{\left(\theta_{i}^{(r)}\right)^{\top}M_{\text{global}}^{(r)}}{\left\|\theta_{i}^{(r)}\right\|\left\|M_{\text{global}}^{(r)}\right\|}$
3:  end for
4:  Apply HDBSCAN to similarity scores
5:  for each cluster $C_{k}$ do
6:     Compute mean cosine similarity $\mu_{k}=\frac{1}{\left\|C_{k}\right\|}\sum_{i\in C_{k}}s_{i}$
7:  end for
8:  Determine benign cluster $\mu_{\text{max}}=\max\left(\{\mu_{k}\}\right)$
9:  Identify benign clients $\mathcal{A}_{\text{benign}}^{(r)}=\left\{i\mid L_{i}=k\text{ and }\mu_{k}=\mu_{\text{max}}\right\}$
10:  Collect benign models ${M}_{\text{benign}}^{(r)}=\left\{\theta_{i}^{(r)}\mid i\in\mathcal{A}_{\text{benign}}^{(r)}\right\}$
11:  return  ${M}_{\text{benign}}^{(r)}$, $\mathcal{A}_{\text{benign}}^{(r)}$
Model Selection. This component refines the set of benign models to select the most representative ones for aggregation. By focusing on models closest to the central tendency, it mitigates the impact of outliers and enhances the robustness of the aggregated model against backdoor attacks.
From the set of benign models ${M}_{\text{benign}}^{(r)}=\left\{\theta_{i}^{(r)}\mid i\in\mathcal{A}_{\text{benign}}^{(r)}\right\}$, we compute the median model parameter vector $\theta_{\text{median}}^{(r)}$. This is done by taking the median of each parameter across all benign models: $\theta_{\text{median}}^{(r)}=\text{median}\left(\left\{\theta_{i}^{(r)}\mid i\in\mathcal{A}_{\text{benign}}^{(r)}\right\}\right),$ where the median is computed element-wise over the parameter vectors.
Next, we calculate the distance between each benign model $\theta_{i}^{(r)}$ and the median model $\theta_{\text{median}}^{(r)}$ using the $L_{1}$ norm: $d_{i}=\left\|\theta_{i}^{(r)}-\theta_{\text{median}}^{(r)}\right\|_{1},\quad\text{for all }i\in\mathcal{A}_{\text{benign}}^{(r)}.$
Models are then ranked based on their distances $d_{i}$. A subset of models with the smallest distances is selected to form the ensemble $\mathcal{E}^{(r)}$: $\mathcal{E}^{(r)}=\left\{\theta_{i}^{(r)}\in{M}_{\text{benign}}^{(r)}\mid d_{i}\leq\epsilon\right\},$ where $\epsilon$ is a statistical measure of the standard deviation of distances from the median to inform the setting of epsilon, ensuring it adapts to the inherent variability in benign updates.
Knowledge Distillation Process. It refines the global model by distilling knowledge from an ensemble of selected benign models, denoted as $\mathcal{E}^{(r)}$, which have been identified through automated clustering and model selection in previous steps. The server utilizes an unlabeled dataset $D_{\text{val}}$, comprising $16\%$ of the total training data, to perform knowledge distillation. Pseudo-labels are generated from the ensemble model outputs to guide the training of the distilled model $M_{\text{distill}}$.
For each sample $x$ in $D_{\text{val}}$, the server computes the logits using each model in the ensemble. Here, logits refer to the raw, unnormalized output scores from the models before applying the softmax function. These logits are averaged to produce the ensemble logits:
[…]
where $f_{M_{i}}(x)$ denotes the logits produced by model $M_{i}$ for input $x$. The pseudo-labels are then generated by applying the softmax function with a temperature parameter $T$ to the averaged logits:
$\tilde{y}(x)=\text{softmax}\left(\frac{\text{Ensemble\_Logits}(x)}{T}\right).$
The temperature parameter $T$ in the softmax function serves as a scaling factor that controls the smoothness of the resulting probability distribution.
The distilled model $M_{\text{distill}}$ is then trained to minimize the Kullback-Leibler (KL) divergence between its output probabilities and the pseudo-labels:
$\mathcal{L}=D_{\text{KL}}\left(\tilde{y}(x)\,\Big{\|}\,\text{softmax}\left(\frac{f_{M_{\text{distill}}}(x)}{T}\right)\right).$
Here, $f_{M_{\text{distill}}}(x)$ represents the logits from the distilled model for input $x$. The KL divergence measures the difference between the two probability distributions, encouraging the distilled model to align closely with the ensemble’s behaviour.
To further stabilize and enhance the generalization of $M_{\text{distill}}$, we employ Stochastic Weight Averaging (SWA) during training. SWA improves generalization by maintaining a running average of the model weights over training epochs, effectively capturing the trajectory of the model parameters as they converge.
During each epoch of the knowledge distillation training, after updating the distilled model $M_{\text{distill}}$ using gradient descent, the SWA model $M_{\text{SWA}}$ is updated as follows:
$M_{\text{SWA}}\leftarrow\frac{n_{\text{SWA}}\cdot M_{\text{SWA}}+M_{\text{distill}}}{n_{\text{SWA}}+1},$
where $n_{\text{SWA}}$ is the number of times the SWA model has been updated. Initially, $M_{\text{SWA}}$ is set to the weights of $M_{\text{distill}}$, and $n_{\text{SWA}}=1$.
Finally, the SWA model $M_{\text{SWA}}$ is set as the updated global model for the next iteration: $M_{\text{global}}^{(r+1)}\leftarrow M_{\text{SWA}}.$
This updated global model integrates the distilled knowledge from the ensemble and the generalization properties provided by SWA. The process of knowledge distillation is in Algorithm 3.
Algorithm 3 Knowledge Distillation Process
0:  $\mathcal{E}^{(r)}$: Ensemble of selected benign models; $D_{\text{val}}$: Unlabeled data for distillation; $T$: Temperature for softmax; $E_{\text{KD}}$: Number of epochs for knowledge distillation; $\eta$: Learning rate
0:  Updated global model $M_{\text{global}}^{(r+1)}$
[…]
2:  Generate pseudo-labels: $\tilde{y}(x)=\text{softmax}\left(\frac{\text{Ensemble\_Logits}(x)}{T}\right)$
3:  for epoch $e=1$ to $E_{\text{KD}}$ do
4:     for each mini-batch $\{x_{b}\}\subset D_{\text{val}}$ do
[…]
6:        Compute loss: $\mathcal{L}=D_{\text{KL}}\left(\tilde{y}(x_{b})\,\Big{\|}\,\text{softmax}\left(\frac{\text{Distill\_Logits}(x_{b})}{T}\right)\right)$
7:        Update $M_{\text{distill}}$ using $SGD$ with $\eta$
8:     end for
9:     Update SWA model: $M_{\text{SWA}}\leftarrow\frac{n_{\text{SWA}}\cdot M_{\text{SWA}}+M_{\text{distill}}}{n_{\text{SWA}}+1}$
10:     $n_{\text{SWA}}\leftarrow n_{\text{SWA}}+1$
11:  end for
12:  Set $M_{\text{global}}^{(r+1)}\leftarrow M_{\text{SWA}}^{(r)}$
13:  return  Updated global model $M_{\text{global}}^{(r+1)}$
V Experiments
We evaluate the effectiveness of the proposed RKD framework in the FL setting under backdoor attack scenarios. We simulate a standard FL environment where multiple clients collaboratively train a global model under the coordination of a central server. The training process is iterative and continues until convergence is achieved.
V-A Datasets and Models
We conducted experiments on three widely used datasets: CIFAR-10, EMNIST, and Fashion-MNIST.
CIFAR-10 [12] consists of 60,000 color images of size $32\times 32$ pixels, categorized into 10 classes. It serves as a comprehensive benchmark for image classification tasks.
EMNIST [6] is an extension of the MNIST dataset, containing $814,255$ handwritten character images across $62$ classes, including digits and letters. The images are grayscale with a $28\times 28$ pixels resolution.
Fashion-MNIST [23] comprises $70,000$ grayscale images of fashion products from $10$ categories, each of size $28\times 28$ pixels, providing a challenging alternative to MNIST dataset.
For each dataset, we employed model architectures suitable for the complexity of the tasks:
For CIFAR-10, we used a ResNet-18 architecture [9], which is well-suited for handling the complexity of colour images and capturing hierarchical features. The model was trained with a batch size of 64 and an initial learning rate of 0.01.
For EMNIST, we utilized a Convolutional Neural Network (CNN) consisting of two convolutional layers, each followed by max pooling and dropout layers to prevent overfitting, and a fully connected layer for classification. The model was trained with a batch size of 64 and a learning rate of 0.001.
For Fashion-MNIST, we implemented a CNN with two convolutional layers, each followed by batch normalization and dropout layers, and a fully connected layer. This architecture aids in normalizing the input features and regularizing the model. The training was conducted with a batch size of 64 and a learning rate of 0.001.
V-B Attack Setup
To evaluate the robustness of the RKD framework against backdoor attacks, we simulated the FL environment with $30$ clients. We considered three scenarios where $20\%$, $40\%$, and $60\%$ of the clients were compromised by an adversary.
Each compromised client injected backdoor triggers into $50\%$ of its local training data. The backdoor trigger was a pattern added to the images, and the labels of the poisoned samples were altered to a target class specified by the adversary. This simulates a realistic attack where malicious clients attempt to implant a backdoor into the global model while maintaining normal performance on clean data.
The compromised clients followed the federated learning protocol but aimed to influence the global model towards the backdoor task. The benign clients trained on their local data without any manipulation.
V-C Attack Methods
To thoroughly evaluate the robustness of our RKD framework, we examined its efficacy against four recent and sophisticated backdoor attack methods. These attacks are designed to circumvent traditional defence mechanisms and present significant challenges in federated learning environments.
Adversarially Adaptive Backdoor Attack to Federated Learning (A3FL)
The A3FL attack [22] enhances backdoor persistence by dynamically adapting the trigger pattern in response to the global training dynamics. Instead of using a static trigger, A3FL continuously optimizes the trigger to remain effective against the evolving global model. It employs an adversarial adaptation loss function and Projected Gradient Descent (PGD) to refine the backdoor trigger. By iteratively updating the trigger based on both the current global model and adversarially crafted models, A3FL ensures that the backdoor remains functional throughout multiple training rounds, making it particularly challenging to detect and mitigate.
Focused-Flip Federated Backdoor Attack (F3BA)
F3BA [7] targets the global model by manipulating a small subset of its parameters with minimal impact on overall performance. The attack calculates the importance of each parameter $w[j]$ concerning the global loss $L_{g}$ using the following metric: $S[j]=-\left(\frac{\partial L_{g}}{\partial w[j]}\right)\cdot w[j]$. Here, $\frac{\partial L_{g}}{\partial w[j]}$ is the gradient of the global loss with respect to parameter $w[j]$, and $\cdot$ denotes element-wise multiplication.
Parameters with the highest importance scores are selected, and their signs are flipped to embed the backdoor. This selective manipulation allows the attacker to implant the backdoor with minimal deviation from normal training behaviour, thereby evading detection.
Distributed Backdoor Attack (DBA)
The DBA [20] spreads a trigger pattern across multiple adversarial clients, enhancing stealth and making detection more difficult. Each compromised client injects a portion of the full trigger into its local training data. When these local models are aggregated, the global model inadvertently learns to associate the combined trigger pattern with the target class. As a result, inputs containing the full trigger pattern are misclassified, effectively executing the backdoor attack without any single client contributing a suspiciously large modification.
Train-and-Scale Backdoor Attack (TSBA)
TSBA [3] aims to evade anomaly detection mechanisms by carefully scaling the model weights after training. The adversarial client first trains its local model with the backdoor trigger until convergence. Then, it scales down the model updates by a factor $\epsilon$ that remains within acceptable bounds defined by the FL protocol. By doing so, the malicious updates appear similar to benign ones in magnitude, thereby bypassing defences that rely on detecting abnormal update sizes. The scaling factor $\epsilon$ is chosen to balance stealth and backdoor effectiveness.
V-D Heterogeneous Setting
In FL, data heterogeneity (Non-IID data distribution) among clients is a common and challenging issue that can significantly impact the learning process and model performance. To simulate realistic federated learning environments with varying degrees of data heterogeneity, we utilize the Dirichlet distribution to partition the datasets among the clients [22].
The Dirichlet distribution allows us to control the level of heterogeneity by adjusting the concentration parameter $\alpha$. Each client $i$ receives a proportion $p_{i,k}$ of data from class $k$, where $p_{i}=[p_{i,1},p_{i,2},\ldots,p_{i,C}]\sim\text{Dirichlet}(\alpha)$ and $C$ is the number of classes. A lower $\alpha$ value leads to more uneven class distributions among clients, simulating higher heterogeneity.
Specifically, we set the $\alpha$ parameter to:
Extreme Heterogeneity: $\alpha=0.5$, $\alpha=0.3$, and $\alpha=0.1$. These lower values of $\alpha$ result in clients having data predominantly from a few classes, leading to highly Non-IID data distributions scenarios.
Moderate Heterogeneity: $\alpha=0.9$ and $\alpha=0.7$. Higher values of $\alpha$ produce more balanced class distributions across clients, representing a moderately heterogeneous setting.
By varying $\alpha$, we create a spectrum of data heterogeneity scenarios to comprehensively evaluate the robustness of the RKD framework under different degrees of Non-IID data distributions. This approach allows us to assess how well the RKD framework can handle challenges posed by data heterogeneity, which is critical for practical FL applications.
Compared Defence Baselines. We evaluate the RKD framework against seven recent FL defences methods: FedAvg [14], FLAME [16], FedDF [13], FedRAD [18], FedBE [5], RLR [17], and FoolsGold (FG) [8].
V-E Evaluation Metrics
We utilized two key evaluation metrics: Main Task Accuracy (MTA) and Attack Success Rate (ASR). These metrics provide a comprehensive understanding of the model’s performance on legitimate tasks and its resistance to backdoor triggers.
Main Task Accuracy (MTA)
MTA measures the classification accuracy of the global model on a clean test dataset $\tau_{m}$, reflecting its ability to correctly predict the true labels of inputs without any backdoor influence. It is defined as: $\text{MTA}=\frac{\left\|\{x\in\tau_{m}\mid f(x)=y\}\right\|}{\left\|\tau_{m}\right\|},$ where $x$ is an input sample from the clean test dataset $\tau_{m}$. $f(x)$ denotes the prediction of the global model $f$ for input $x$. $y$ is the true label corresponding to $x$. ||$\cdot$|| ⋅ | represents the cardinality of the set.
A higher MTA indicates that the model performs well on the primary classification task, correctly identifying inputs as per their true labels.
Attack Success Rate (ASR)
ASR evaluates the effectiveness of the backdoor attack by measuring the proportion of poisoned inputs that the global model misclassifies into the attacker’s target class. It is calculated on a test dataset containing backdoor triggers $\tau_{b}$: $\text{ASR}=\frac{\left\|\{x\in\tau_{b}\mid f(x)=y_{\text{target}}\}\right\|}{\left\|\tau_{b}\right\|},$ where $x$ is an input sample from the poisoned test dataset $\tau_{b}$ that contains backdoor triggers. $f(x)$ is the prediction of the global model $f$ for input $x$. $y_{\text{target}}$ is the target label that the attacker intends the model to output for backdoor inputs.
A lower ASR signifies better robustness against backdoor attacks, as it indicates that the model is less likely to misclassify backdoor inputs into the attacker’s target class.
The goal of an effective Defence mechanism like the RKD framework is to maintain a high MTA while minimizing the ASR. This balance ensures that the model retains its performance on legitimate data while being resilient to manipulation attempts by adversaries. In our experiments, we focus on achieving this balance to demonstrate the RKD framework’s capability to defend against sophisticated backdoor attacks without degrading the overall model performance.
V-F Experimental Results
We evaluated the robustness of the RKD framework against advanced backdoor attacks in FL. The models were trained under non-IID data distributions, measuring the MTA and ASR. To ensure reliability, all experiments were repeated five times with different data resampling, with confidence intervals reported at a significance level of $\rho=0.01$.
V-F1 Defence Against A3FL Attack
Under highly heterogeneous non-IID conditions ($\alpha=0.3$), RKD demonstrated significant resilience against the A3FL attack on the CIFAR-10 and Fashion-MNIST datasets. As illustrated in Figures 1 and 2, RKD substantially reduced the ASR while maintaining high MTA compared to baseline methods.
(a) $20\%$ of A3FL
(b) $40\%$ of A3FL
(c) $60\%$ of A3FL
Figure 1: Performance of baselines and RKD on CIFAR-10 under Non-IID ($\alpha=0.3$), evaluated against $20\%$, $40\%$, and $60\%$ A3FL attacker clients.
(a) $20\%$ of A3FL
(b) $40\%$ of A3FL
(c) $60\%$ of A3FL
Figure 2: Performance of baselines and RKD on Fashion-MNIST under Non-IID ($\alpha=0.3$), evaluated against $20\%$, $40\%$, and $60\%$ A3FL attacker clients.
The superior performance of RKD against A3FL can be attributed to the synergy of its methodological components:
First, RKD employs cosine similarity to assess the directional alignment of model updates with the global model. RKD effectively detects malicious updates that embed backdoor triggers without significantly altering parameter norms—a limitation in methods like FLAME that cluster based on parameter magnitudes. This approach enables RKD to capture angular deviations introduced by adversarial models, isolating backdoor updates that might otherwise appear benign in terms of update size. Specifically, as detailed in our Automated Clustering methodology, the server computes the cosine similarity between each client’s model parameters and the global parameter vector. This metric captures significant directional deviations characteristic of malicious behaviour, allowing for effective detection.
Second, the use of HDBSCAN for clustering allows RKD to identify outliers without requiring a predefined number of clusters. By clustering the similarity scores obtained from cosine similarity, rather than the high-dimensional parameter vectors, RKD reduces computational complexity and avoids the curse of dimensionality. The dynamic adjustment of HDBSCAN’s minimum cluster size parameter, as described in our methodology, enables RKD to adapt to evolving model updates during training. This adaptability enhances clustering robustness, ensuring the accurate grouping of benign models while isolating potential malicious ones.
Third, after identifying and excluding malicious models through clustering, RKD utilizes the Model Selection component to further refine the aggregation process. By computing the median model parameters within the benign cluster and selecting models closest to the median model, RKD reduces the influence of any remaining outliers that may not have been filtered out during clustering. This selection is based on the distances of models from the median model parameters. As outlined in our methodology, this step mitigates the impact of potential undetected malicious updates and enhances the robustness of the aggregated model against backdoor attacks.
Fourth, the Knowledge Distillation Process plays a crucial role in consolidating the insights from the selected benign models. By distilling knowledge from the ensemble of these models into a global model using a temperature-scaled softmax function and the Kullback-Leibler divergence loss, RKD aligns the distilled model closely with the consensus of benign models. This process, as detailed in our methodology, mitigates the risk of incorporating hidden backdoor triggers that might not have been detected during clustering, ensuring that the global model retains high performance on legitimate tasks.
Integrating SWA during the distillation process further refines the global model by smoothing the training trajectory. SWA averages the weights of the model over multiple epochs, leading to convergence towards flatter minima in the loss landscape. This contributes to better generalization and strengthens the defence against backdoor attacks by reducing the model’s sensitivity to small perturbations introduced by adversarial updates. As per our methodology, SWA enhances the stability and generalization capability of the distilled model, ensuring robustness against backdoor attacks and maintaining high performance across Non-IID data distributions.
Crucially, RKD restricting access to the global model by disseminating the aggregated global model only to clients identified as benign. By withholding the updated global model from adversarial clients, RKD prevents attackers from adapting their strategies based on the latest model parameters. This measure is particularly effective against adaptive attacks like A3FL, which rely on feedback from the global model to refine their backdoor triggers. As described in our methodology, this approach limits the influence of malicious clients and enhances the overall security of the federated learning process.
By sending a minimally perturbed version of the global model (with a noise scale of $1\times 10^{-4}$) to suspected malicious clients, we obscured the exact state of the model, effectively preventing attackers from adapting their strategies. We evaluated the Perturbed Global Model (PGM) strategy within the RKD framework to maintain model performance while mitigating adaptive attack risks and reducing the likelihood of attackers realizing they have been identified. Comparing RKD with PGM (denoted RKD(PGM)) to RKD without PGM (denoted RKD) under various attack scenarios of A3FL, we found that RKD(PGM) maintained an average accuracy closely matching that of RKD, indicating that the minimal perturbations did not adversely affect performance.
V-F2 Defense Against F3BA Attack
RKD effectively defends against F3BA on CIFAR-10 and EMNIST datasets under non-IID conditions ($\alpha=0.5$), as substantiated by Figures 3 and 4. Using cosine similarity-based clustering, RKD filters out anomalies from compromised clients and integrates knowledge distillation to maintain low ASR and high accuracy.
RKD’s iterative training enhances the global model’s accuracy and resilience, demonstrating its superiority over methods like FedAvg. This is particularly evident in high attacker ratios of $40\%$ and $60\%$, highlighting RKD’s robust defence against sophisticated attacks like F3BA.
(a) $20\%$ of F3BA
(b) $40\%$ of F3BA
(c) $60\%$ of F3BA
Figure 3: Performance of baselines and RKD on CIFAR-10 under Non-IID ($\alpha=0.5$), evaluated against $20\%$, $40\%$, and $60\%$ F3BA attacker clients.
(a) $20\%$ of F3BA
(b) $40\%$ of F3BA
(c) $60\%$ of F3BA
Figure 4: Performance of baselines and RKD on EMNIST under Non-IID ($\alpha=0.5$), evaluated against $20\%$, $40\%$, and $60\%$ F3BA attacker clients.
V-F3 Defense Against DBA Attack
RKD effectively defends against the Distributed Backdoor Attack (DBA) on CIFAR-10 and EMNIST under non-IID settings ($\alpha=0.9$), as shown in Figures 5 and 6. Using cosine similarity-based clustering, RKD detects and isolates malicious updates. Median model selection ensures that only benign models contribute to the global model, minimizing subtle backdoor triggers.
During knowledge distillation, RKD synthesizes insights from vetted models into a robust aggregated model, enhancing generalizability and security. Compared to methods like FedDF, FedRAD, and FedBE, RKD provides superior protection by meticulously analyzing and distilling knowledge from selected models. This enables RKD to maintain high accuracy while significantly reducing the ASR, demonstrating its effectiveness against sophisticated attacks like DBA.
(a) $20\%$ of DBA
(b) $40\%$ of DBA
(c) $60\%$ of DBA
Figure 5: Performance of baselines and RKD on CIFAR-10 under Non-IID ($\alpha=0.9$), evaluated against $20\%$, $40\%$, and $60\%$ DBA attacker clients.
(a) $20\%$ of DBA
(b) $40\%$ of DBA
(c) $60\%$ of DBA
Figure 6: Performance of baselines and RKD on EMNIST under Non-IID ($\alpha=0.9$), evaluated against $20\%$, $40\%$, and $60\%$ DBA attacker clients.
V-F4 Defense Against TSBA Attack
The RKD framework robustly defends against TSBA on CIFAR-10 and EMNIST under Non-IID conditions ($\alpha=0.5$), as shown in Figures 7 and 8. RKD detects and mitigates TSBA manipulations, maintaining high accuracy and low ASR even with increased attacker ratios. Unlike other methods that falter under poisoned conditions, RKD excels with clean and poisoned datasets.
(a) $20\%$ of TSBA
(b) $40\%$ of TSBA
(c) $60\%$ of TSBA
Figure 7: Performance of baselines and RKD on CIFAR-10 under Non-IID ($\alpha=0.5$), evaluated against $20\%$, $40\%$, and $60\%$ TSBA attacker clients.
(a) $20\%$ of TSBA
(b) $40\%$ of TSBA
(c) $60\%$ of TSBA
Figure 8: Performance of baselines and RKD on EMNIST under Non-IID ($\alpha=0.5$), evaluated against $20\%$, $40\%$, and $60\%$ TSBA attacker clients.
The Impact of Heterogeneous Degree. We evaluated baseline defence methods and the RKD framework under varying Non-IID levels, including moderate ($\alpha=0.7$) and extreme ($\alpha=0.3,0.1$), as shown in Figure 9.
Baseline methods, despite high accuracy on clean inputs, falter in detecting embedded trigger patterns, leading to elevated ASR. In contrast, RKD consistently maintains robust defence, effectively countering the heightened threats of extreme heterogeneity. This highlights RKD’s superior adaptability and effectiveness where baseline defences struggle. While extreme heterogeneity impacts accuracy, RKD outperforms baseline methods by excluding backdoor models and ensuring robust performance and enhanced defence against backdoor attacks.
Refer to captionRefer to captionRefer to caption
(a) Non-IID=$0.7$
(b) Non-IID=$0.3$
(c) Non-IID=$0.1$
Figure 9: Performance impact of heterogeneous degrees on baselines and RKD on Fashion-MNIST, evaluated against $60\%$ F3BA attacker clients.
Empirical Analysis of $Q$ Sensitivity
To evaluate the impact of the minimum cluster size ($Q$) on MTA and ASR, we performed experiments using the CIFAR-10 dataset under a Non-IID setting with $30$ clients, of which $40\%$ were malicious and executing A3FL backdoor attacks. The results are summarized in Table I.
TABLE I: Impact of $Q$ on MTA and ASR
$Q$ MTA ASR
Fixed $Q=2$ High High
Observations: Small cluster size allows malicious updates to dominate clusters, maintaining MTA but elevating ASR.
Fixed $Q=20$ Decreased, especially in later rounds Low
Observations: Larger clusters exclude many benign updates, reducing ASR. Over multiple rounds, this exclusion accumulates, leading to a global model that lacks sufficient diversity from benign updates, which negatively impacts the model’s ability to generalize and perform well on the main task.
Dynamic $Q$ Consistently high Low
Observations: Effectively excludes malicious updates while retaining benign ones, ensuring high MTA and reduced ASR.
Scalability. RKD enhances scalability by applying cosine similarity to model updates before clustering, transforming high-dimensional parameter vectors into scalar similarity scores. This dimensionality reduction significantly lowers computational complexity, making the clustering process more efficient. By avoiding clustering in the high-dimensional parameter space, RKD reduces both the time and resources required for defence operations. As shown in Table II, RKD’s defence time is $42.029$ seconds, substantially faster than FedDF and FedBE, which require $141.714$ and $198.765$ seconds, respectively. While RLR exhibits the shortest defence time, it compromises on detection accuracy due to its simplistic approach. FLAME is slightly more efficient in defence time, but RKD achieves a better balance between performance and robustness. These results highlight RKD’s overall efficiency and scalability, demonstrating that its methodological design—specifically the use of cosine similarity and efficient clustering—provides robust defence without incurring significant computational overhead.
Method Defense Time (seconds)
RLR 0.020
FLAME 37.064
RKD 42.029
FedDF 141.714
FedBE 198.765
TABLE II: Comparative Defense Times
V-F5 Ablation Study
We conducted ablation studies to evaluate the effectiveness of each component within the RKD framework against sophisticated backdoor attacks. Specifically, we analyzed the impact of removing key components: Automated Clustering, Model Selection, and Knowledge Distillation, as shown in Figure 10.
RKD without Clustering and Model Selection. Removing both the Automated Clustering and Model Selection components, and leaving only the knowledge distillation process, significantly weakens the framework’s defenses. While it performs reasonably well under a $20\%$ DBA attack, as the proportion of adversarial clients rises to $40\%$, its defensive capabilities drop sharply. At a $60\%$ F3BA attacker ratio, the model suffers from severe performance degradation, with misclassifications aligning with attackers’ objectives. This highlights that without clustering, RKD is unable to effectively identify and isolate malicious updates, allowing backdoor attacks to compromise the global model.
RKD without Model Selection. Excluding the model selection component while retaining HDBSCAN clustering, this variant effectively manages a $20\%$ DBA attack ratio. However, with a $40\%$ attack ratio, performance is noticeably decreased. Under a $60\%$ F3BA attacker ratio, the removal of model selection further impairs RKD’s defense, resulting in increased ASR and reduced MTA. Even after the clustering phase effectively isolates most malicious updates, some malicious or anomalous updates may still be present. Without the median, these residual outliers can disproportionately influence the aggregated model, leading to instability in its parameters when evaluated under attack success rate. These findings underscore the critical role of model selection in refining the aggregation process by selecting models closest to the benign cluster centroid, thus enhancing robustness against higher proportions of attackers.
RKD without Knowledge Distillation. Retaining both clustering and model selection but omitting the Knowledge Distillation module, this variant successfully identifies and excludes malicious models, achieving a lower ASR. However, it suffers from a significant drop in accuracy, approximately $20\%$ lower than the full RKD framework. This underscores the crucial role of KD in transferring knowledge from the selected ensemble of models to the global model, which enhances both accuracy and generalization, particularly across Non-IID settings. Without KD, the framework is unable to effectively integrate and refine the knowledge from ensemble models, leading to poor performance in Non-IID environments despite successful isolation of malicious updates. KD specifically addresses the challenge of generalization in Non-IID data scenarios, ensuring the model remains robust and effective.
(a) $20\%$ of DBA
(b) $40\%$ of DBA
(c) $60\%$ of F3BA
Figure 10: Ablation study of RKD method against attacks.
V-G Additional Experiments
Defense Against A3FL Attack Under IID Conditions. As shown in Figure 11, RKD outperforms baseline models in defending against the A3FL attack on the CIFAR-10 dataset under IID conditions, across varying attacker ratios ($20\%$, $40\%$, and $60\%$). RKD’s robustness against adversarial attacks in IID settings underscores its effectiveness as a defence mechanism. The baseline models also demonstrate some inherent robustness at lower attacker ratios ($20\%$ and $40\%$), which can be attributed to the uniform data distribution in IID settings that allows models to learn more generalizable patterns, making it harder for attacks to succeed.
(a) $20\%$ of A3FL
(b) $40\%$ of A3FL
(c) $60\%$ of A3FL
Figure 11: Performance of baselines and RKD on CIFAR-10 under IID settings against A3FL attacker clients.
VI Conclusion and Future Work
We present RKD, a novel framework designed to mitigate sophisticated backdoor attacks in FL under Non-IID conditions. Our extensive evaluations on CIFAR-10, EMNIST, and Fashion-MNIST datasets demonstrated RKD’s effectiveness in maintaining high accuracy levels while significantly reducing attack success rates. RKD’s automated clustering, model selection, and knowledge distillation components collectively enhance the robustness and integrity of FL systems. Our results underscore RKD’s superiority over existing defence methods, offering a robust solution to preserve the efficacy and security of FL in diverse and adversarial environments.
Acknowledgments
We gratefully acknowledge the support of the Saudi Arabian Ministry of Education, the Saudi Arabian Cultural Bureau in London, and Umm Al-Qura University in Makkah. We are thankful to the High-End Computing facility at Lancaster University for providing essential computational resources. We also acknowledge the scholarship awarded by the 3rd IEEE Conference on Secure and Trustworthy Machine Learning, which facilitated our participation. Finally, we express our sincere gratitude to Dr. Matheus Aparecido do Carmo Alves for his invaluable review.
-A Threat Model 1 - Malicious Clients Have Access to Benign Updates.
This scenario assumes an extremely informed adversary with capabilities far exceeding the typical assumptions in RKD threat models. Specifically, we consider the case where malicious clients have complete access to all benign updates in each round. While this represents an unrealistic adversarial scenario in federated learning, it serves as a valuable stress test to evaluate the robustness of RKD. Below, we outline the adversary coordination strategy: Adversary 1: Leveraging its knowledge of all benign updates, Adversary 1 selects its parameter update such that the new average, denoted as avg_new, is shifted to a specific point in the parameter space aligned with the adversarial objective. This manipulation ensures that avg_new reflects the desired direction of the attack while remaining plausible within the model update space. Adversary 2: Adversary 2 submits an update identical to avg_new. Since the clustering mechanism in RKD relies on cosine similarity values, Adversary 2’s update achieves a perfect cosine similarity score of 1 with avg_new, making it indistinguishable from benign updates.
Experimental Evidence.
Our experiments, illustrated in Figure 12, confirm that RKD’s reliance on cosine similarity clustering, while effective against most attacks, can be insufficient against such a highly coordinated strategy. Adversary 2’s inclusion in the benign cluster allows the adversaries to significantly influence the global model aggregation, causing a systematic shift in the global model’s parameters toward their objectives.
-B Threat Model 2 - Malicious Clients Do Not Have Access to Benign Updates.
This practical scenario assumes that malicious clients have only partial knowledge, meaning they lack direct access to benign updates and must approximate them indirectly. Adversary Strategy: In the Observation Phase, malicious clients behave benignly for multiple rounds, submitting updates that closely follow the global model to observe its evolution. In the Approximation Phase, they use these observations to estimate the average benign update, allowing them to generate updates that align with the attack strategy outlined in Threat Model 1.
Challenges for Adversarial Approximation:
Heterogeneity of Benign Updates. In FL, benign updates exhibit significant variability due to the non-IID nature of client data. This variability introduces inconsistencies that cannot be inferred from observing aggregated global updates alone. Additionally, the global model updates provide an oversimplified view of the collective contributions, masking the nuanced behavior of individual benign clients and limiting the accuracy of adversarial estimations.
RKD’s Clustering Based on Cosine Similarity. RKD clusters client updates using cosine similarity, emphasizing the directional alignment of each update with the global model. Even if adversaries attempt to approximate the average benign updates, achieving precise directional alignment across all malicious clients is inherently challenging. Malicious clients performing backdoor attacks typically have diverse and conflicting objectives, resulting in updates that do not uniformly align with the global model’s optimization direction. This variability makes it difficult for malicious clients to consistently approximate the benign update direction, preventing them from seamlessly blending into the benign clusters.
Moreover, HDBSCAN dynamically adjusts clustering boundaries based on the evolving distribution of similarity scores each round. This continuous adaptation makes it difficult for adversaries to predict the exact clustering outcomes.
RKD’s Knowledge Distillation Adds Randomness. RKD employs knowledge distillation using pseudo-labels generated by the ensemble on a validation dataset. This process introduces additional randomness into the final global model, disrupting adversarial efforts to align with benign updates.
Potential Impact on RKD.
The success of this attack hinges on the adversary’s ability to accurately estimate benign updates and replicate their alignment with the global model, thereby infiltrating benign clusters or evading detection.
Experimental Evidence.
Our experiments, illustrated in Figure 12, validate that RKD effectively defends against Threat Model 2. Despite malicious clients attempting to approximate benign updates, their approximations consistently deviate from actual benign updates. Even when the observation phase is prolonged, providing adversaries with additional iterations to gather data, their approximations remain inefficient and misaligned.
The accuracy trends of RKD and FedAvg under Threat Model 2, where adversaries lack direct access to benign updates and attempt to approximate them indirectly. This scenario simulates a more practical and realistic adversarial condition in federated learning. The results underscore the robustness of RKD and its superior performance in practical threat scenarios. At the same time, traditional aggregation methods like FedAvg in federated learning, lacking robust defense mechanisms, are unable to counteract even partially informed adversaries, resulting in a dramatic accuracy loss that underscores the necessity of designing effective defenses to ensure trustworthy global model aggregation in FL.
(a) Threat Model 1
(b) Threat Model 2
Figure 12: Impact of Threat Model 1 and 2 on Model Accuracy under RKD and FedAvg.
-C The Anti-Distillation Backdoor Attack (ADBA.)
The ADBA [26], initially designed for centralized machine learning scenarios utilizing knowledge distillation, has been adapted in this work to the federated learning paradigm. Originally, ADBA embeds a robust backdoor into a teacher model by dynamically optimizing a trigger and using a shadow model to ensure backdoor transfer through distillation. In extending this attack to FL, the global model is treated as the teacher, while client models act as distributed learners, effectively mirroring the hierarchical structure of knowledge distillation. Compromised clients inject backdoors by optimizing a crafted trigger using projected gradient descent while adhering to FL constraints. The results presented in 13 for RKD, in comparison with FedAvg and other baseline methods, underscore RKD’s effectiveness in detecting and mitigating ADBA backdoor attacks.
(a) $20\%$ of ADBA
(b) $40\%$ of ADBA
(c) $60\%$ of ADBA
Figure 13: Performance of baselines and RKD on CIFAR-10 under Non-IID settings against ADBA attacker clients.
-D Performance Under Non-IID and IID Conditions Without Attacks
We evaluated the performance of RKD and baseline methods under both Non-IID and IID data distributions in the absence of attacks. The results, illustrated in Figure 14, demonstrate that the attack success rate remains close to zero across all methods when no attacks are present. Additionally, models trained under IID conditions consistently achieve higher accuracy compared to those trained under Non-IID.
This difference is attributed to the uniform distribution of data in IID settings, which facilitates more effective training and generalization. Conversely, the lower accuracy observed in Non-IID models reflects the challenges posed by data heterogeneity, highlighting the importance of robust training strategies to handle Non-IID data distributions effectively.
(a) Non-IID
(b) IID
Figure 14: Performance of baselines and RKD on CIFAR-10 under Non-IID and IID conditions with no attack.
-E Defense Against F3BA Attack Under IID Conditions.
Figure 15 illustrates RKD’s effective defense against the F3BA attack on the CIFAR-10 dataset under IID conditions. RKD consistently achieves lower ASR and higher MTA compared to baseline methods across different attacker ratios. This superior performance highlights RKD’s potential as a reliable defense in adversarial federated learning environments. Baseline models show some robustness against $20\%$ of F3BA attackers, likely due to the IID settings enabling the models to learn generalizable patterns that are less susceptible to manipulation by adversarial clients.
(a) $20\%$ of F3BA
(b) $40\%$ of F3BA
(c) $60\%$ of F3BA
Figure 15: Performance of baselines and RKD on CIFAR-10 under IID settings against F3BA attacker clients.